The following facts pertain to a lease agreement between Leasing Company and Brown Company, a lessee. (Round all numbers to the nearest cent.)

 Inception date: March 1, 2019

 Annual lease payment due at the beginning of each year, beginning with March 1, 2019: €13,604.79.

 Unguaranteed residual value after 6 years is €5,000.00.  Lease term: 6 years

 Economic life of leased equipment: 10 years

 Lessor’s cost: €53,000.00; fair value of asset at March 1, 2019, €68,000.00

 Lessor’s implicit rate: 10%; lessee’s incremental borrowing rate 10% Instructions

1. Discuss the nature of this lease to Leasing Company. 2. Prepare a lease amortization schedule for Brown Company for the 6-year lease term. 3. Prepare the journal entries on the lessee’s books to reflect the signing of the lease agreement and to record the payments and expenses related to this lease for the years 2019 and 2020. Brown’s annual accounting period ends on December 31.
Answer #1

Answer:-

wn PV of periodic rental payments = 13604.79*PV of Annuity due of 1 for n =6 ,i-10%=13604.79*4.79079 =65178
PV of guaranteed residual value = 5000 * PV of 1 for n=6,i-10%=5000*0.56447=2822   
Lease receivable at inception = 65178+2822 = 68000
1 For Brown company it is a capital Lease as the present value of minimum lease payments exceeds 90% of the fair value of assets.
For leasing company it is a sales type lease as the fair value of assets exceeds the lessor's cost.
2 Year Annual Lease payment Interest Reduction of Lease Liability Lease Liability
3/1/2019 68000
3/1/2019 13604.79 13604.79 54395.21
3/1/2020 13604.79 5439.52 8165.27 46229.94
3/1/2021 13604.79 4622.99 8981.8 37248.15
3/1/2022 13604.79 3724.81 9879.98 27368.17
3/1/2023 13604.79 2736.82 10867.97 16500.2
3/1/2024 13604.79 1650.02 11954.77 4545.43
4/30/2024 5000 454.57 4545.43 0
Total 86628.74 18628.74 68000
Date Account Title Debit-$ Credit-$
1-Mar-19 Lease Equipment 68000
Lease Liability 68000
to Record Lease
Lease Liability 13604.79
Cash 13604.79
To record Lease payment
12/31/2019 Interest Expense 4532.93 5439.52/12*10
Interest payable 4532.93
To record interest
3/1/2020 Lease Liability 8165.27
Interest payable 4532.93
Interest expense 906.59 5439.52-4532.93
Cash 13604.79
To record Lease payment
12/31/2020 Interest Expense 3852.49 4622.99/12*10
Interest payable 3852.49
To record interest

if you have any query please ask me in comment box i am here to helps you don't give direct Thumbs down.if you satisfied my work give Thumbs UP

******THANK YOU*******